OpenStudy (anonymous):

What should I do when the variables in my equations don't cancel out when I add them together? I'm trying to use the elimination method!

OpenStudy (anonymous):

4x + 3y = -1 3x + y = 3 is the problem, but y and 3y don't cancel out??

Directrix (directrix):

Which variable did you try to eliminate?

OpenStudy (anonymous):

I want to eliminate 3y?

Directrix (directrix):

4x + 3y = -1 3x + y = 3 ----> Multiply this equation by -3 --> -3x -3y = -3 4x + 3y = -1 -3x -3y = -3 ------------ x + 0y = -4 x = -4 ---> value of x @CharlesDavid --> Solve for y using either of the original equations.

OpenStudy (anonymous):

Oh ok, I think I get it now. Thanks that really helps! I know how to solve for y, so thanks for all your help c:

Directrix (directrix):

What did you get for y?